# On the Characteristics of the Conjugate Function Enabling Effective Dual Decomposition Methods

### v1

@@ III FGOR Characteristic of the Dual Function
Let us now investigate the FGOR characteristic of the dual function $g$ for problem (2) that is directly inherited from the conjugate function $f^{*}$, cf.  § II.
Let $\bm{\lambda}\in{\rm I\!R}^{m}$ be the Lagrange multiplier associated to the equality constraint of problem (2). Then the dual function is given by
@@
From (7), it is clear that the dual function $g$ is based on a restriction of $-f^{*}$ to a linear space. Moreover, from basic calculus rules, we have
@@
Consider minimizing the function $f_{2}$ in Example 1 with an equality constraint $\mathbf{Ay}=\mathbf{0}$. The corresponding optimization problem is given below.
@@
[…]
We are now ready to show our main result about the FGOR characteristic of the dual function $g$, cf.  (7).
@@
The proof is a direct consequence of Proposition 1 and Assumption 3. More specifically, from Assumption 3, there exists a ray $\mathcal{R}\in\mathcal{A}$ such that $\mathcal{R}\subseteq R(\mathbf{A}^{\mbox{\scriptsize T}})$. Let =$\mathcal{R}~{}{=\{\bm{\nu}+\alpha\bm{\eta}\ |\ \alpha\geq 0\}}$ for some $\bm{\nu},\bm{\eta}\in{\rm I\!R}^{n}$. Now consider the inverse image $\mathcal{I}_{\mathcal{R}}$ of $\mathcal{R}$ under the linear mapping $\mathbf{A}^{\mbox{\scriptsize T}}$. In particular, we have
=$\mathcal{I}_{\mathcal{R}}=\{\bm{\kappa}\ |\ \mathbf{A}^{\mbox{\scriptsize T}}\bm{\kappa}\in\mathcal{R}\}=\{\bm{\lambda}+\alpha\bm{\mu}\ |\ \alpha\geq 0\},$ (12)
@@
We describe in the following example two instances of dual functions associated to problem (11) to demonstrate the assertions of Corollary 1.
@@
[…]
Finally, let us remark that the conjugate function and the dual function are both gradient-Lipschitz continuous. Assumption 1 is sufficient to yield such Lipschitzian properties, although we won’t take this up in our subsequent stepsize designs. We note that the aforementioned Lipschitzian properties are more general than the existing results. Details are deferred to Appendix B for completeness.
@@ IV On the Application of FGOR on Dual Subgradient Algorithm
where $\gamma_{k}>0$ is the stepsize, $\mathbf{s}^{(k)}$ is a subgradient of $-g$ at $\bm{\lambda}^{(k)}\in{\rm I\!R}^{m}$, and $k$ represents the iteration index.
@@
We now give some insight into how the stepsize $\gamma_{k}$ values should be chosen by considering problem (11) and its associated dual function $g_{1}$ depicted in Figure 2b. The idea is very simple. Loosely speaking the stepsize rule is to use a constant stepsize as long as $\lambda^{(k)}$ is in a region of FGOR [cf.  $\mathcal{F}_{1}=\mathcal{F}_{11}\cup\mathcal{F}_{12}$, Figure 2b], making the subgradient method yield a constant increase in the dual objective function in every iteration. For instance, moving from $\lambda^{(0)}$ to $\lambda^{(1)}$, from $\lambda^{(1)}$ to $\lambda^{(2)}$, and so forth is due to constant stepsizes since $\lambda^{(0)},\lambda^{(1)},\lambda^{(2)}\in\mathcal{F}_{11}$, which in turn makes the subgradient method able to have constant increments, cf.  Figure 2b. The important point is to make the subgradient method (14) exit the region $\mathcal{F}_{1}$ effectively and reach quickly the region ${\mathcal{F}}_{1}^{\star}$ in which the dual solution $\lambda_{1}^{\star}$ resides. As soon as the subgradient method exits the region $\mathcal{F}_{1}$, the stepsize can be restored to any desirable existing one to guarantee convergence. Note that the constant stepsize rule we propose can precede existing ones by furnishing potential advantages as we will further explain in subsequent sections.
@@ IV-A Formalizing the FGOR on Dual Subgradient
for some $\beta>0$. Figure 3 illustrates the idea of the condition by depicting two rays $\{\bm{\lambda}+\alpha\bm{\mu}\ |\ \alpha\geq 0\}$ and $\{\tilde{\bm{\lambda}}+\alpha\tilde{\bm{\mu}}\ |\ \alpha\geq 0\}$. Note that the gradient $\nabla g$ is a constant vector over each ray. However, the condition (15) is satisfied only over the first ray.
$\bm{\lambda}$ $\bm{\mu}$ $\bm{\bar{\lambda}}$ $\bm{\bar{\mu}}$
$\nabla g$ $\nabla g$ $\bm{\lambda}$ $\bm{\mu}$ $\bm{\bar{\lambda}}$ $\bm{\bar{\mu}}$
Figure 3: An illustration of condition (15): the condition is satisfied only over the ray $\{\bm{\lambda}+\alpha\bm{\mu}\ |\ \alpha\geq 0\}$. LABEL:sub@subfig:grad-g-parallel-to-line Level curves of a dual function $g~{}{:{\rm I\!R}^{3}\to{\rm I\!R}}$ and two rays with the FGOR characteristic. LABEL:sub@subfig:grad-g-not-parallel-to-line-plane-view Corresponding plan view.
@@
IV-B Determining a Ray with the FGOR Characteristic
In the case of (15), the main issue to be settled is the determination of a ray $\{\bm{\lambda}+\alpha\bm{\mu}\ |\ \alpha\geq 0\}$ over which the condition (15) is affirmative. In this respect, the following equivalence plays an important role:
$\nabla g(\bm{\lambda}+\alpha\bm{\mu})=-\beta{\bm{\mu}}$
$\iff-\mathbf{A}\nabla f^{*}\big{(}\mathbf{A}^{\mbox{\scriptsize T}}(\bm{\lambda}+\alpha\bm{\mu})\big{)}=-\beta{\bm{\mu}}-\mathbf{b}$ (18)
$\iff-\mathbf{A}\nabla f^{*}(\bm{\nu}+\alpha\bm{\eta})=-\beta(\mathbf{A}\mathbf{A}^{\mbox{\scriptsize T}})^{-1}\mathbf{A}{\bm{\eta}}-\mathbf{b},$ (19)
where (18) follows from (8), and (19) follows from that $\bm{\lambda}$ and $\bm{\mu}$ are inverse images of some $\bm{\nu}$ and $\bm{\eta}$ under the linear mapping $\mathbf{A}^{\mbox{\scriptsize T}}$, cf.  (12). Note that $\{\bm{\nu}+\alpha\bm{\eta}\ |\ \alpha\geq 0\}$ is the image of the ray $\{\bm{\lambda}+\alpha\bm{\mu}\ |\ \alpha\geq 0\}$ under $\mathbf{A}^{\mbox{\scriptsize T}}$. The equivalence (19) confirms that a ray $\{\bm{\lambda}+\alpha\bm{\mu}\ |\ \alpha\geq 0\}$ over which condition (15) is affirmative can be determined by seeking for a ray =$\mathcal{R}=\{\bm{\nu}+\alpha\bm{\eta}\ |\ \alpha\geq 0\}\in\mathcal{A}$ such that
@@ IV-B Determining a Ray with the FGOR Characteristic
$\mathcal{R}\subseteq R(\mathbf{A}^{\mbox{\scriptsize T}}).$ (21)
We start by noting that the subdifferential $\partial f(\cdot)$ evaluated at $\texttt{bnd}~{}\mathcal{Y}$ [cf.  Remark 3] and the collection of rays $\mathcal{A}$ over which the gradients of the conjugate function are constant [cf.  Proposition 1, (6), (10)] are intimately connected. Consequently, for fix $\bar{\mathbf{y}}\in\texttt{bnd}~{}\mathcal{Y}$, together with the inversion rule for subgradient relations [42, Prop. 11.3], we have
$\{\bm{\nu}+\alpha\bm{\eta}\ |\ \alpha\geq 0\}\subseteq\partial f(\bar{\mathbf{y}})\iff\forall~{}\alpha\geq 0,\ \bar{\mathbf{y}}=\nabla f^{*}(\bm{\nu}+\alpha\bm{\eta}).$ (22)
The equivalence (22) suggests that if a ray of the form $\{\bm{\nu}+\alpha\bm{\eta}\ |\ \alpha\geq 0\}$ satisfies $\bar{\mathbf{y}}=\nabla f^{*}(\bm{\nu}+\alpha\bm{\eta})$, then the ray resides in $\partial f(\bar{\mathbf{y}})$. Thus, in line with this observation, together with condition (20) and condition (21), we cast the following convex feasibility problem toward yielding the ray with the intended characteristics:
minimize0subject $\begin{array}[]{ll}\mbox{minimize}&0\\ \mbox{subject to}&\mathbf{A}\bar{\mathbf{y}}-\mathbf{b}=\beta(\mathbf{A}\mathbf{A}^{\mbox{\scriptsize T}})^{-1}\mathbf{A}{\bm{\eta}}\\ &\{\bm{\nu}+\alpha\bm{\eta}\ |\ \alpha\geq 0\}\subseteq\partial f(\bar{\mathbf{y}})\\ &\{\bm{\nu}+\alpha\bm{\eta}\ |\ \alpha\geq 0\}\subseteq R(\mathbf{A}^{\mbox{\scriptsize T}}),\end{array}$ (23)
where the variables are ${\bm{\nu}}$ and ${\bm{\eta}}$. It is straightforward to see that problem (23) can equivalently be formulated as
minimize0subject $\begin{array}[]{ll}\mbox{minimize}&0\\ \mbox{subject to}&\mathbf{A}\bar{\mathbf{y}}-\mathbf{b}=\beta(\mathbf{A}\mathbf{A}^{\mbox{\scriptsize T}})^{-1}\mathbf{A}{\bm{\eta}}\\ &{\bm{\nu}}\in\partial f_{0}(\bar{\mathbf{y}})+\partial\delta_{\mathcal{Y}}(\bar{\mathbf{y}})\\ &{\bm{\eta}}\in\partial\delta_{\mathcal{Y}}(\bar{\mathbf{y}})\\ &\mathbf{A}^{\mbox{\scriptsize T}}\bm{\lambda}={\bm{\nu}}\\ &\mathbf{A}^{\mbox{\scriptsize T}}\bm{\mu}={\bm{\eta}},\end{array}$ (24)
where the variables are ${\bm{\nu}}$, ${\bm{\eta}}$, ${\bm{\lambda}}$, and ${\bm{\mu}}$. If problem (24) is solved with solution $({\bm{\nu}}^{\star},{\bm{\eta}}^{\star},{\bm{\lambda}}^{\star},{\bm{\mu}}^{\star})$, the ray we seek is simply given by
$\mathcal{R}^{\star}(\bar{\mathbf{y}})=\{\bm{\lambda}^{\star}+\alpha\bm{\mu}^{\star}\ |\ \alpha\geq 0\}.$ (25)
An illustration is given in Figure 4LABEL:sub@subfig:feasible-problem1-1. If the problem is infeasible, one may choose a different $\bar{\mathbf{y}}\in\texttt{bnd}~{}\mathcal{Y}$.
@@
In practice, one should have available the knowledge of $\partial f_{0}(\bar{\mathbf{y}})$ to solve problem (24). It is worth noting that, in a distributed optimization setting, the global objective function $f_{0}$ is often based on local objective functions of the involved subsystems. Consequently, $\partial f_{0}(\bar{\mathbf{y}})$, unlike other problem data, cannot be determined based on only the subsystems’ local information. Therefore, this issue has to be dealt with through coordination among the subsystems. If such coordination is to be avoided, an alternative is to relax problem (24) by dropping the second constraint that entails the main coupling among subsystems. Consequently, the fourth constraint $\mathbf{A}^{\mbox{\scriptsize T}}\bm{\lambda}={\bm{\nu}}$ becomes obsolete and the resulting relaxed formulation is given by
@@
minimize0subject $\begin{array}[]{ll}\mbox{minimize}&0\\ \mbox{subject to}&\mathbf{A}\bar{\mathbf{y}}-\mathbf{b}=\beta(\mathbf{A}\mathbf{A}^{\mbox{\scriptsize T}})^{-1}\mathbf{A}{\bm{\eta}}\\ &\mathbf{A}^{\mbox{\scriptsize T}}\bm{\mu}={\bm{\eta}}\\ &{\bm{\eta}}\in\partial\delta_{\mathcal{Y}}(\bar{\mathbf{y}}),\end{array}$ (26)
where the variables are ${\bm{\eta}}$ and ${\bm{\mu}}$. A careful examination of problem (26) shows that it can be solved in two steps: 1) solve first two equations to uniquely yield the solutions $\tilde{\bm{\mu}}=(1/\beta)(\mathbf{A}\mathbf{\bar{y}}-\mathbf{b})$ and $\tilde{\bm{\eta}}=\mathbf{A}^{\mbox{\scriptsize T}}\tilde{\bm{\mu}}$, 2) check whether the condition $\tilde{\bm{\eta}}\in\partial\delta_{\mathcal{Y}}(\bar{\mathbf{y}})$ holds. If the condition is affirmative, then the unique solution of problem  (26) is $(\tilde{\bm{\eta}},\tilde{\bm{\mu}})$. Otherwise, a different $\bar{\mathbf{y}}\in\texttt{bnd}~{}\mathcal{Y}$ may be chosen and the process is repeated.
When considering problem (26) instead of (24), there are trade-offs indeed. Unlike (24), problem (26) can be solved independently by each subsystem. This is because it is quite common in practice that the problem parameters $\mathbf{A}$ and $\mathcal{Y}$ of problem (1) are known by subsystems as far as distributed optimization settings are concerned (e.g., consensus problem). On the other hand, Unlike (24), the solution to problem (26) doesn’t directly yield a ray as sought. Nevertheless, if feasible, we let $\tilde{\bm{\lambda}}=\tilde{\bm{\mu}}$ and consider the line given by
$\mathcal{L}^{\star}(\bar{\mathbf{y}})=\{\tilde{\bm{\lambda}}+\alpha\tilde{\bm{\mu}}\ |\ \alpha\in{\rm I\!R}\}.$ (27)
It is easily seen that $\mathcal{L}^{\star}(\bar{\mathbf{y}})$ contains a ray of the intended FGOR characteristic, cf.  Figure 4LABEL:sub@subfig:feasible-problem1-2. This suggests the choice of the initialization point $\bm{\lambda}^{(0)}$ from $\mathcal{L}^{\star}(\bar{\mathbf{y}})$.

### v3

@@ III FGOR Characteristic of the Dual Function
The FGOR characteristic of the dual function $g$ for problem (2) is directly inherited from the conjugate function $f^{*}$, cf.  § II. To this end, we start by letting $\bm{\lambda}\in{\rm I\!R}^{m}$ be the Lagrange multiplier associated to the equality constraint of problem (2). Then the dual function is given by
@@
From (10), it is clear that the dual function $g$ is based on a restriction of $f^{*}$ to a linear space. Moreover, from basic calculus rules, we have
@@
Consider minimizing $f_{2}$ in Example 1 with an equality constraint $\mathbf{Ay}=\mathbf{0}$. Then the corresponding optimization problem is
@@
The sets $R(\mathbf{A}_{1}^{\mbox{\scriptsize T}})$ and $R(\mathbf{A}_{2}^{\mbox{\scriptsize T}})$, where $\mathbf{A}_{1}=[-5\ 3]$ and $\mathbf{A}_{2}=[1\ 1]$ are depicted in Figure2LABEL:sub@subfig:example-dual-function-restrictions. Moreover, the rays $\mathcal{R}(\mathbf{v}_{1})$ and $\mathcal{R}(\mathbf{v}_{2})$ [cf.  (12)], where $\mathbf{v}_{1}=[-3\ 1.8]^{\mbox{\scriptsize T}}$ and $\mathbf{v}_{2}=[3.2\ 3.2]^{\mbox{\scriptsize T}}$, are also shown in the same figure. Clearly, $\mathcal{R}(\mathbf{v}_{1})\subseteq R(\mathbf{A}_{1}^{\mbox{\scriptsize T}})$ and $\mathcal{R}(\mathbf{v}_{2})\subseteq R(\mathbf{A}_{2}^{\mbox{\scriptsize T}})$, respectively. Thus Assumption 3 holds for problem (14).
We note that Assumption 3 is inherently problem-dependent and may not hold in all cases. The lack of a general verification criterion for this assumption constitutes a caveat of the proposed approach, which must be addressed separately using problem-specific ingenuity. We are now ready to show our main result about the FGOR characteristic of the dual function $g$, cf.  (10).
@@
The proof is a direct consequence of Proposition 1, Assumption 3, and $\mathbf{A}^{\mbox{\scriptsize T}}\in{\rm I\!R}^{n\times m}$ with $m<n$ and $\texttt{rank}~\mathbf{A}=m$. More specifically, from Assumption 3, there exists a ray $\mathcal{R}\in\mathcal{A}$ such that $\mathcal{R}\subseteq R(\mathbf{A}^{\mbox{\scriptsize T}})$. Let $\mathcal{R}~{=\{\bm{\nu}+\alpha\bm{\eta}\ |\ \alpha\geq 0\}}$ for some $\bm{\nu},\bm{\eta}\in{\rm I\!R}^{n}$. Now consider the inverse image $\mathcal{I}_{\mathcal{R}}$ of $\mathcal{R}$ under the linear mapping $\mathbf{A}^{\mbox{\scriptsize T}}$. In particular, we have
$\mathcal{I}_{\mathcal{R}}=\{\bm{\kappa}\ |\ \mathbf{A}^{\mbox{\scriptsize T}}\bm{\kappa}\in\mathcal{R}\}=\{\bm{\lambda}+\alpha\bm{\mu}\ |\ \alpha\geq 0\},$ (15)
@@
In the following example, we describe two instances of dual functions associated to problem (14) to demonstrate the assertions of Corollary 1.
@@
Figure 2LABEL:sub@subfig:example-dual-functions depicts the dual functions $g_{1}$ and $g_{2}$ of problem (14) when $\mathbf{A}=\mathbf{A}_{1}=[-5\ 3]$ and $\mathbf{A}=\mathbf{A}_{2}=[1\ 1]$, respectively. The functions $g_{1}$ and $g_{2}$ are based on restrictions of $f^{*}$ to $R(\mathbf{A}_{1}^{\mbox{\scriptsize T}})$ and $R(\mathbf{A}_{2}^{\mbox{\scriptsize T}})$, respectively, cf.  Figure 2LABEL:sub@subfig:example-dual-function-restrictions. The figure shows both $g_{1}$ and $g_{2}$ have regions with constant gradients.
Finally, let us remark that the conjugate function and the dual function are both gradient-Lipschitz continuous. Assumption 1 is sufficient to yield such Lipschitzian properties, although we won’t take this up in our subsequent stepsize designs. We note that the aforementioned Lipschitzian properties are more general than the existing results. Details are deferred to Appendix F for completeness.
@@ IV On the Application of FGOR on Dual Subgradient Algorithm
where $\gamma_{k}>0$ is the stepsize, $\mathbf{s}^{(k)}$ is a subgradient of g-g at $\bm{\lambda}^{(k)}\in{\rm I\!R}^{m}$, and $k$ represents the iteration index.
@@
We now give some insight into how the stepsize $\gamma_{k}$ values should be chosen by considering problem (14) and its associated dual function $g_{1}$ depicted in Figure 2b. The idea is very simple. Roughly speaking, the stepsize rule is to use a constant stepsize as long as $\lambda^{(k)}$ is in a region of FGOR [cf.  $\mathcal{F}_{1}=\mathcal{F}_{11}\cup\mathcal{F}_{12}$, Figure 2b], making the subgradient method yield a constant increase in the dual objective function in every iteration. For instance, moving from $\lambda^{(0)}$ to $\lambda^{(1)}$, from $\lambda^{(1)}$ to $\lambda^{(2)}$, and so forth is due to constant stepsizes since $\lambda^{(0)},\lambda^{(1)},\lambda^{(2)}\in\mathcal{F}_{11}$, which in turn makes the subgradient method able to have constant increments, cf.  Figure 2b. The important point is to make the subgradient method (17) exit the region $\mathcal{F}_{1}$ effectively and reach quickly the region ${\mathcal{F}}_{1}^{\star}$ in which the dual solution $\lambda_{1}^{\star}$ resides. As soon as the subgradient method exits the region $\mathcal{F}_{1}$, the stepsize can be restored to any desirable existing one to guarantee convergence. Note that the constant stepsize rule we propose can precede existing ones by furnishing potential advantages as we will further explain in subsequent sections.
@@ IV-A Formalizing the FGOR on Dual Subgradient
for some $\beta>0$. In other words, besides being a constant vector, condition (18) requires that the gradient $\nabla g$, when evaluated along the ray, aligns with the ray itself. We refer to the set of all rays that satisfy condition (18) as the restricted-FGOR (RFGOR) region. Figure 3 illustrates the idea of the condition by depicting two rays $\{\bm{\lambda}_{1}+\alpha\bm{\mu}_{1}\ |\ \alpha\geq 0\}$ and $\{\bm{\lambda}_{2}+\alpha\bm{\mu}_{2}\ |\ \alpha\geq 0\}$. Note that the gradient $\nabla g$ is a constant vector over each ray. However, the condition (18) is satisfied only over the first ray.
$\bm{\lambda}_{1}$ $\bm{\mu}_{1}$ $\bm{\lambda}_{2}$ $\bm{\mu}_{2}$
$\nabla g$ $\nabla g$ $\bm{\lambda}_{1}$ $\bm{\mu}_{1}$ $\bm{\lambda}_{2}$ $\bm{\mu}_{2}$
Figure 3: An illustration of condition (18): the condition is satisfied only over the ray $\{\bm{\lambda}_{1}+\alpha\bm{\mu}_{1}\ |\ \alpha\geq 0\}$. LABEL:sub@subfig:grad-g-parallel-to-line Level curves of a dual function $g~{:{\rm I\!R}^{3}\to{\rm I\!R}}$ and two rays with the FGOR characteristic. LABEL:sub@subfig:grad-g-not-parallel-to-line-plane-view Corresponding plan view.
@@
IV-B Determining a Ray Ensuring (18)
The computation of a ray in the form $\{\bm{\lambda}+\alpha\bm{\mu}\ |\ \alpha\geq 0\}$, over which condition  (18) holds, is carried out through a properly formulated feasibility problem, as detailed in the sequel. Let us first note the following equivalent conditions for (18):
$\forall\alpha\geq 0,~\nabla g(\bm{\lambda}+\alpha\bm{\mu})=-\beta{\bm{\mu}}$
$\iff\forall\alpha\geq 0,~-\mathbf{A}\nabla f^{*}\big(\mathbf{A}^{\mbox{\scriptsize T}}(\bm{\lambda}+\alpha\bm{\mu})\big)=-\beta{\bm{\mu}}-\mathbf{b}$ (21)
$\iff\forall\alpha\geq 0,~-\mathbf{A}\nabla f^{*}(\bm{\nu}+\alpha\bm{\eta})=-\beta(\mathbf{A}\mathbf{A}^{\mbox{\scriptsize T}})^{-1}\mathbf{A}{\bm{\eta}}-\mathbf{b},$ (22)
where (21) follows from (11) and (22) follows from that $\bm{\lambda}$ and $\bm{\mu}$ are inverse images of some $\bm{\nu}$ and $\bm{\eta}$ under the linear mapping $\mathbf{A}^{\mbox{\scriptsize T}}$ [cf.  (15)]. Consequently, the feasibility of the optimization problem
$\begin{array}[]{ll}\mbox{minimize}&0\\ \mbox{subject to}&\forall\alpha\geq 0,~-\mathbf{A}\nabla f^{*}(\bm{\nu}+\alpha\bm{\eta})\\ &\hskip 60.00009pt=-\beta(\mathbf{A}\mathbf{A}^{\mbox{\scriptsize T}})^{-1}\mathbf{A}{\bm{\eta}}-\mathbf{b}\\ &\mathbf{A}^{\mbox{\scriptsize T}}\bm{\lambda}=\bm{\nu}\\ &\mathbf{A}^{\mbox{\scriptsize T}}\bm{\mu}=\bm{\eta}\end{array}$ (23)
with variables $\bm{\lambda}$}, $\bm{\mu}$}, $\bm{\nu}$}, and $\bm{\eta}$}, is sufficient for (18).
At this point, we rely on Proposition 1 to yield a more constrained formulation of (23). More specifically, Proposition 1 claims that there are rays in ${\rm I\!R}^{n}\setminus\texttt{int}~\mathcal{V}$ over which the gradient $\nabla f^{*}$ is constant and is characterized by some $\bar{\mathbf{y}}\in\texttt{bnd}~\mathcal{Y}$ [cf.  (9)]. Thus, in line with this claim, let us impose an additional constraint $\forall~\alpha\geq 0,\ \bar{\mathbf{y}}=\nabla f^{*}(\bm{\nu}+\alpha\bm{\eta})$, where $\bar{\mathbf{y}}\in\texttt{bnd}~\mathcal{Y}$. As a result, we have the following problem, which is more constrained than (23):
$\begin{array}[]{ll}\mbox{minimize}&0\\ \mbox{subject to}&-\mathbf{A}\bar{\bm{y}}=-\beta(\mathbf{A}\mathbf{A}^{\mbox{\scriptsize T}})^{-1}\mathbf{A}{\bm{\eta}}-\mathbf{b}\\ &\forall\alpha\geq 0,~\nabla f^{*}(\bm{\nu}+\alpha\bm{\eta})=\bar{\bm{y}}\\ &\mathbf{A}^{\mbox{\scriptsize T}}\bm{\lambda}=\bm{\nu}\\ &\mathbf{A}^{\mbox{\scriptsize T}}\bm{\mu}=\bm{\eta},\end{array}$ (24)
where the variables are $\bm{\lambda}$}, $\bm{\mu}$}, $\bm{\nu}$}, and $\bm{\eta}$}. By using [Rockafellar-98, Prop. 11.3], (24) can equivalently be reformulated as
$\begin{array}[]{ll}\mbox{minimize}&0\\ \mbox{subject to}&-\mathbf{A}\bar{\bm{y}}=-\beta(\mathbf{A}\mathbf{A}^{\mbox{\scriptsize T}})^{-1}\mathbf{A}{\bm{\eta}}-\mathbf{b}\\ &\{\bm{\nu}+\alpha\bm{\eta}\ |\ \alpha\geq 0\}\subseteq\partial f(\bar{\mathbf{y}})\\ &\mathbf{A}^{\mbox{\scriptsize T}}\bm{\lambda}=\bm{\nu}\\ &\mathbf{A}^{\mbox{\scriptsize T}}\bm{\mu}=\bm{\eta},\end{array}$ (25)
where the second constraint of (24) is updated accordingly. It is straightforward to see that the second constraint in (25) can break into two constraints ${\bm{\nu}}\in\partial f_{0}(\bar{\mathbf{y}})+N_{\mathcal{Y}}(\bar{\mathbf{y}})$ and ${\bm{\eta}}\in N_{\mathcal{Y}}(\bar{\mathbf{y}})$ [cf. proof of Proposition 1, (7), (8)]. Thus, problem (25) is equivalent to
$\begin{array}[]{ll}\mbox{minimize}&0\\ \mbox{subject to}&\mathbf{A}\bar{\mathbf{y}}-\mathbf{b}=\beta(\mathbf{A}\mathbf{A}^{\mbox{\scriptsize T}})^{-1}\mathbf{A}{\bm{\eta}}\\ &{\bm{\nu}}\in\partial f_{0}(\bar{\mathbf{y}})+N_{\mathcal{Y}}(\bar{\mathbf{y}})\\ &{\bm{\eta}}\in N_{\mathcal{Y}}(\bar{\mathbf{y}})\\ &\mathbf{A}^{\mbox{\scriptsize T}}\bm{\lambda}={\bm{\nu}}\\ &\mathbf{A}^{\mbox{\scriptsize T}}\bm{\mu}={\bm{\eta}},\end{array}$ (26)
where the variables are $\bm{\nu}$}, $\bm{\eta}$}, $\bm{\lambda}$}, and $\bm{\mu}$}. It is worth noting that
@@ IV-B Determining a Ray Ensuring (18)
${\mathcal{R}}^{\star}(\bar{\mathbf{y}})=\{\hat{\bm{\lambda}}+\alpha\hat{\bm{\mu}}\ |\ \alpha\geq 0\}.$ (27)
An illustration is given in Figure 4LABEL:sub@subfig:feasible-problem1-1. If problem (26) is infeasible, one may choose a different $\bar{\mathbf{y}}\in\texttt{bnd}~\mathcal{Y}$.
Figure 4: An illustration of a ray ${\mathcal{R}}^{\star}(\mathbf{\bar{y}})$ [cf.  (27)] associated to problem (14), where $\mathbf{A}=[-0.05\ 0.03]$, $\mathbf{b}=\mathbf{0}$, $\mathbf{\bar{y}}=[-2\ 2]^{\text{T}}$, and $\beta=1$. The sets $N_{\mathcal{Y}_{2}}(\mathbf{\bar{y}})$ and $\partial f_{2}(\mathbf{\bar{y}})+N_{\mathcal{Y}_{2}}(\mathbf{\bar{y}})$ are depicted by the regions shaded in pink and green, respectively. LABEL:sub@subfig:feasible-problem1-1 ${\mathcal{R}}^{\star}(\mathbf{\bar{y}})$ given by solving (26): $\hat{\bm{\nu}}=[-3.25\ 1.95]^{\text{T}}$, $\hat{\bm{\eta}}=[-0.008\ 0.0048]^{\text{T}}$, $\hat{\bm{\lambda}}=65$, $\hat{\bm{\mu}}=0.16$. LABEL:sub@subfig:feasible-problem1-2 ${\mathcal{R}}^{\star}(\mathbf{\bar{y}})$ given by solving (28): $\bm{\tilde{\eta}}=[-0.008\ 0.0048]^{\text{T}}$$\bm{\tilde{\mu}}=0.16$. Then $\hat{\bm{\nu}}=c\bm{\tilde{\eta}}$, $\hat{\bm{\eta}}=\bm{\tilde{\eta}}$, $\hat{\bm{\lambda}}=c\bm{\tilde{\mu}}$, $\hat{\bm{\mu}}=\bm{\tilde{\mu}}$, where $c=500$, cf.  (29).
@@
In practice, one should have available the knowledge of $\partial f_{0}(\bar{\mathbf{y}})$ to solve problem (26). Note that in a distributed optimization setting, the global objective function $f_{0}$ is often based on local objective functions of the involved subsystems. Consequently, $\partial f_{0}(\bar{\mathbf{y}})$, unlike other problem data, cannot be determined based on only the subsystems’ local information. Therefore, this issue has to be dealt with through coordination among the subsystems. If such coordination is to be avoided, an alternative is to relax problem (26) by dropping the second constraint that entails the main coupling among subsystems. Consequently, the fourth constraint $\mathbf{A}^{\mbox{\scriptsize T}}\bm{\lambda}={\bm{\nu}}$ becomes obsolete and the resulting relaxed formulation is given by
@@
$\begin{array}[]{ll}\mbox{minimize}&0\\ \mbox{subject to}&\mathbf{A}\bar{\mathbf{y}}-\mathbf{b}=\beta(\mathbf{A}\mathbf{A}^{\mbox{\scriptsize T}})^{-1}\mathbf{A}{\bm{\eta}}\\ &{\bm{\eta}}\in N_{\mathcal{Y}}(\bar{\mathbf{y}})\\ &\mathbf{A}^{\mbox{\scriptsize T}}\bm{\mu}={\bm{\eta}},\end{array}$ (28)
[…]
$\hskip-5.69054pt\hat{\bm{\mu}}=\bm{\tilde{\mu}},\ \hat{\bm{\eta}}=\bm{\tilde{\eta}},\ \hat{\bm{\nu}}=c\bm{\tilde{\eta}},\ \hat{\bm{\lambda}}=c(\mathbf{AA}^{\mbox{\scriptsize T}})^{-1}\mathbf{A}\bm{\tilde{\eta}}=c\bm{\tilde{\mu}}$ (29)
[…]
Note that problem (28) can be solved independently by each subsystem, unlike (26). This is because it is quite common in practice that the parameters $\mathbf{A}$ and $\mathcal{Y}$ of problem (1) are known by subsystems as far as distributed optimization settings are concerned (e.g., consensus problem). Moreover, it is worth noting that both problems (28) and (26) do not require knowing in advance whether Assumption 3 is satisfied. To see this, let us first remark the following:
Remark 3
The feasibility of problem (28) [or (26)] confirms the satisfaction of Assumption 3.
Proof:
The feasibility of (28) [or (26)] ensures the existence of an RFGOR ray $\mathcal{R}^{\star}(\mathbf{\bar{y}})\subseteq\texttt{dom}~g$, cf.  (27),(29). Consequently, $\nabla f^{*}$, when evaluated over the ray $\mathbf{A}^{\mbox{\scriptsize T}}\mathcal{R}^{\star}(\mathbf{\bar{y}})$ [which is trivially in ${R}(\mathbf{A}^{\mbox{\scriptsize T}})$], becomes a constant vector [cf.  (11)], yielding the FGOR characteristic. ∎
As such, even without explicitly knowing whether Assumption 3 holds, one can simply solve problem (28) and compute an RFGOR ray $\mathcal{R}^{\star}(\mathbf{\bar{y}})$, if feasible. Then we provide the initialization point $\bm{\lambda}^{(0)}$ from $\mathcal{R}^{\star}(\mathbf{\bar{y}})$ for the dual subgradient method. Let us next discuss the corresponding algorithm.